# Enhanced Textual Feature Extraction for Visual Question Answering: A Simple Convolutional Approach

### v2

@@ 1 Introduction
Visual Question Answering (VQA) [1] has emerged as an increasingly interesting area of research at the intersection of computer vision and natural language processing. The objective of VQA is to answer questions about a given image, requiring models capable of understanding both visual and textual modalities. One particular task in the VQA domain that merits additional attention is the counting task. This specific task requires the model to quantitatively identify the number of objects in a given image. Despite its apparent simplicity, this task poses significant challenges, as it demands precise identification, localization, and counting of objects, all while avoiding potential pitfalls like multiple counting or not acknowledging occluded items.
Advanced sequential models, such as Transformers, have seen staggering success in various tasks and have been widely adopted. However, these models are not always the best choice for every task, particularly in settings where balancing model performance with computational cost is essential. As highlighted in previous studies that evaluate models under specific conditions [2], carefully considering model selection and exploring simple yet effective improvements is also crucial for optimizing task performance.
In this paper, we pose the question: Are complex sequential models the most suitable approach for handling textual modality in VQA tasks, especially on the original VQA-v2 dataset? To investigate this problem, we conduct rigorous experiments applying both complex models like Transformer Encoder and attention-based models, as well as simpler structures such as RNNs and CNNs on the VQA-v2 dataset. We focus our analysis on different text feature extractions and their impact on model accuracy. We discover that simpler models, specifically those good at capturing local interdependencies within the text, can provide improvements over some complex models. Furthermore, the incorporation of convolutional layers into simpler structures like Gated Recurrent Units (GRU), forming what we term as ConvGRU, proves quite feasible.
Our main contributions are threefold: (1) We present ConvGRU, a GRU-based text encoder enhanced by utilizing convolutional layers. Experimental results demonstrate that leveraging local textual features with ConvGRU can marginally improve VQA accuracy without increasing computational costs. (2) We conduct an extensive comparison of various VQA text encoders, revealing that complex models such as Transformer Encoders may underperform compared to simpler architectures better at extracting local textual features.
@@ 3.1 Problem Formulation
The primary objective of Visual Question Answering (VQA) is to accurately predict an answer given an image and a corresponding natural-language question. Therefore, the VQA task can be conceptualized as a classification challenge, where the goal is to identify the most probable answer $\hat{a}$ from a predefined set of possible answers, based on the provided image $i$ and a question $q$. Formally, this is expressed as:
@@ 3.2 Method Overview
Our methodology seeks to improve VQA task outcomes by focusing on enhanced text processing techniques, leveraging the foundational work of [22]. We introduce a model that primarily innovates in textual feature extraction through a novel Convolutional GRU framework, while also efficiently integrating pre-existing advanced methods for image feature extraction and a counting component to address both visual data and numerical queries effectively.
At the heart of our proposed model, as depicted in Figure 1, is the integration of a Convolutional GRU framework for improved extraction of textual features from questions. For image modality, we employ a pre-trained R-CNN model for extracting features, complemented by a stacked attention mechanism and a counting module from the base model. This combination offers a comprehensive solution for the VQA challenge.
@@ 3.3 Standard Models
This section outlines the established models and methods integrated into our VQA framework. Our choice to utilize well-documented techniques from previous research underscores our textual modality innovations and supports comparative analysis with other textual approaches.
@@ 3.4 Convolutional GRU for Text Encoding
We propose a one-dimensional convolutional approach as a more effective alternative for feature extraction. By employing kernels of sizes 2 and 3, we capture bi-gram and tri-gram features, respectively, which are then processed through distinct padding strategies-asymmetrical for bi-grams to prioritize the first word of the sequence, and symmetrical for tri-grams to preserve context evenly. Similar to the observations in [25] regarding the shift problem encountered with even-numbered convolutions in image processing, we identify a parallel issue in the realm of one-dimensional text convolutions. To avoid the potential complexity and overfitting that might arise from directly addressing this shift problem, we choose to simply add a unit of padding at the beginning of sequences to accurately identify the initial words of a question, which are crucial for defining its category. For instance, "what" typically signals a query about an object, while "what color" suggests a question about color. The utility of this asymmetric head padding technique is also validated by the results in Section 4.3.
@@ 4.1 Dataset and Evaluation Metric
VQA-v2 Dataset. A notable change of Visual Question Answering was the transition from the VQA dataset to VQA-v2 dataset in order to address a number of biases and language priors found in the original dataset [3]. For instance, "blue" is frequently selected as the answer to questions beginning with "What color is …" regardless of the image content, which may lead to improper correlations between the question text and the answer. To cope with these constraints, VQA-v2 introduces complementary pairs, wherein there is a "complementary" image that yields a different answer to the same question for every question linked to an image. In our experiments, VQA-v2 dataset is chosen as our primary focused task, and we assess our model on validation set after training on the training set.
@@ 4.3 Results
Table 2: Comparison of different textual models. "SA" and "MHA" refer to Self-Attention and Multi-Head Self-Attention respectively, and all configurations are the same except for textual encoders.
@@ 4.5 How Short Can VQA Questions Get?
As shown in Figure 3, we use bar chart to represent the distribution, with the X-axis indicating the length of questions and the Y-axis indicating the frequency of questions at each length. In our comprehensive analysis of 658,111 questions from VQA-v2 dataset, we found that a significant majority, 636,894 questions (96.78%), have lengths ranging from 3 to 10 words. Notably, 569,437 questions (86.53%) are within the 4 to 8-word range. A steep decrease in question frequency is observed with increasing length. Focusing on the "How many" counting questions (72,012 in total), a predominant 61,516 questions (85.42%) fall within the 5 to 8-word range, with questions exceeding 10 words represent a mere 4.1% of counting queries.
These insights highlight the predominance of short questions in VQA-v2 dataset, demonstrating the effectiveness of smaller convolution kernels for local feature extraction. The prevalence of question-type determinants at the start of queries ("How many", "Is there", "Do", etc.) further advocates for the superiority of compact kernels. This observation also aligns with the performance dip for k=4 scenarios depicted in Table 3, reinforcing smaller kernels’ utility in enhancing textual modality within VQA frameworks.
@@ 4.6 Qualitative Results
  • As shown in Case (a), the original GRU incorrectly predicts the answer to be "blue", likely confusing the color of the background with that of the text. Conversely, both the ConvGRU(3) and the ConvGRU(2+3) models accurately identify the correct answer as "white". This illustrates the capability of these enhanced models to capture key phrase pairs such as "color of" and "wording on" through 2-grams, and richer context through 3-grams like "the color of" and "of the wording," which effectively utilize the convolutional layers to discern the finer details of the question.
  • In Case (b), the GRU model gives a completely wrong prediction of "picture" as it relies solely on simplistic word-level features. The ConvGRU(3) model improves by predicting "sitting"—a rationally deduced yet contextually incorrect response, revealing that 3-gram features effectively capture a person’s state or action in a specific location. The ConvGRU(2+3), by incorporating both 2-gram and 3-gram features, correctly comprehends the query’s intent and the subject’s context, leading to an accurate "reading" prediction. This case exemplifies how combining different n-grams can enrich linguistic feature understanding and enhance alignment with image features for more accurate predictions.
  • In both the counting task and the yes/no task, as shown in Case (c) and Case (d), ConvGRU(2+3) showcases best performance. The ability to understand relational context and specific details allows it to correctly count "3" boys on the field and accurately confirm the presence of avocado on the sandwich with a "yes". These outcomes illustrate the effectiveness of integrating multi-gram features into GRUs.
@@ 5 Conclusion
This paper explored the effectiveness of complex versus simpler models for capturing textual features in Visual Question Answering (VQA) tasks, specifically addressing whether complex sequential models are optimal for this purpose. Our experiments demonstrated that the ConvGRU model, designed to capture local text features, marginally outperformed the baseline GRU model, while advanced models like Self-Attention and Transformer Encoders showed reduced effectiveness. The primary advantage of ConvGRU lies in its ability to enhance performance with minimal computational cost, and it supports the importance of balancing model complexity with efficiency.

### v3

@@ 1 Introduction
Visual Question Answering (VQA) [1] has emerged as an increasingly interesting area of research at the intersection of computer vision and natural language processing. The objective of VQA is to answer questions about a given image, requiring models capable of understanding both visual and textual modalities. One particular task in the VQA domain that attracts additional attention is the counting task. This specific task requires the model to quantitatively identify the number of objects in a given image. Despite its apparent simplicity, this task demands precise identification, localization, and counting of objects, and avoiding potential pitfalls like multiple counting or not acknowledging occluded items.
Advanced sequential models, such as Transformers, have seen great success in various tasks and have been widely adopted. However, these models are not always the best choice for every task, particularly in scenarios where balancing model performance with computational cost is essential. As highlighted in previous studies that evaluate models under specific conditions [2], carefully considering model selection and exploring simple yet effective improvements is also crucial for optimizing task performance.
In this paper, we pose the question: Are complex sequential models always the most suitable approach for handling textual modality in VQA tasks? To investigate this problem, we conduct comprehensive experiments applying both complex models like Transformer Encoder and attention-based models, as well as simpler structures such as RNNs and CNNs on a well-established VQA-v2 dataset. We focus our analysis on different text feature extraction methods and their impact on model accuracy. We discover that simpler models, specifically those good at capturing local interdependencies within the text, can provide improvements over some complex models. Furthermore, the incorporation of convolutional layers into simpler structures like Gated Recurrent Units (GRU), forming what we term as ConvGRU, proves quite feasible.
Our main contributions are threefold: (1) We present ConvGRU, a GRU-based text encoder enhanced by applying convolutional layers. Experimental results demonstrate that leveraging local textual features with ConvGRU can marginally improve VQA accuracy without increasing computational costs. (2) We conduct an extensive comparison of various VQA text encoders, and find that complex models such as Transformer Encoders may underperform compared to simpler architectures better at extracting local textual features.
@@ 3.1 Problem Formulation
The primary objective of Visual Question Answering (VQA) is to accurately predict an answer given an image and a corresponding natural-language question. Therefore, the VQA task can be regarded as a classification problem, and its goal is to identify the most probable answer $\hat{a}$ from a predefined set of possible answers, based on the provided image $i$ and a question $q$. Formally, this is expressed as:
@@ 3.2 Method Overview
Our methodology seeks to improve VQA accuracies by focusing on enhanced text processing techniques, leveraging the foundational work of [22]. We introduce a model that primarily innovates in textual feature extraction through a novel Convolutional GRU framework, and also integrate existing methods for image feature extraction and a counting component to address both visual data and numerical queries effectively.
At the heart of our proposed model, as depicted in Figure 1, is the integration of a Convolutional GRU framework for improved extraction of textual features from questions. For image modality, we employ a pre-trained R-CNN model for extracting features, accompanied by a stacked attention mechanism and a counting module from the base model.
@@ 3.3 Standard Models
We directly adopt several established methods to highlight our contributions in the textual modality and give a fair comparison of different textual encoders under the same setting.
@@ 3.4 Convolutional GRU for Text Encoding
We propose a one-dimensional convolutional approach as a more effective alternative for feature extraction. By employing kernels of sizes 2 and 3, we capture bi-gram and tri-gram features, respectively, which are then processed through distinct padding strategies-asymmetrical for bi-grams to prioritize the first word of the sequence, and symmetrical for tri-grams to preserve context evenly. Similar to the observations in [25] regarding the shift problem encountered with even-numbered convolutions in image processing, we identify a parallel issue in the realm of one-dimensional text convolutions. To avoid the potential complexity and overfitting that might arise from directly addressing this shift problem, we choose to simply add a unit of padding at the beginning of sequences to accurately identify the initial words of a question, which are crucial for defining its category. For instance, “what” typically signals a query about an object, while “what color” suggests a question about color. The utility of this asymmetric head padding technique is also validated by the results in Section 4.3.
@@ 4.1 Dataset and Evaluation Metric
VQA-v2 Dataset. A notable change of Visual Question Answering was the transition from the VQA dataset to VQA-v2 dataset in order to address a number of biases and language priors found in the original dataset [3]. For instance, “blue” is frequently selected as the answer to questions beginning with “What color is …” regardless of the image content, which may lead to improper correlations between the question text and the answer. To cope with these constraints, VQA-v2 introduces complementary pairs, wherein there is a “complementary” image that yields a different answer to the same question for every question linked to an image. In our experiments, VQA-v2 dataset is chosen as our primary focused task, and we assess our model on validation set after training on the training set.
@@ 4.3 Results
Table 2: Comparison of different textual models. “SA” and “MHA” refer to Self-Attention and Multi-Head Self-Attention respectively, and all configurations are the same except for textual encoders.
@@ 4.5 How Short Can VQA Questions Get?
As shown in Figure 3, we use bar chart to represent the distribution, with the X-axis indicating the length of questions and the Y-axis indicating the frequency of questions at each length. In our comprehensive analysis of 658,111 questions from VQA-v2 dataset, we found that a significant majority, 636,894 questions (96.78%), have lengths ranging from 3 to 10 words. Notably, 569,437 questions (86.53%) are within the 4 to 8-word range. A steep decrease in question frequency is observed with increasing length. Focusing on the “How many” counting questions (72,012 in total), a predominant 61,516 questions (85.42%) fall within the 5 to 8-word range, with questions exceeding 10 words represent a mere 4.1% of counting queries.
These insights highlight the predominance of short questions in VQA-v2 dataset, demonstrating the effectiveness of smaller convolution kernels for local feature extraction. The prevalence of question-type determinants at the start of queries (“How many”, “Is there”, “Do”, etc.) further advocates for the superiority of compact kernels. This observation also aligns with the performance dip for k=4 scenarios depicted in Table 3, reinforcing smaller kernels’ utility in enhancing textual modality within VQA frameworks.
@@ 4.6 Qualitative Results
  • As shown in Case (a), the original GRU incorrectly predicts the answer to be “blue”, likely confusing the color of the background with that of the text. Conversely, both the ConvGRU(3) and the ConvGRU(2+3) models accurately identify the correct answer as “white”. This illustrates the capability of these enhanced models to capture key phrase pairs such as “color of” and “wording on” through 2-grams, and richer context through 3-grams like “the color of” and “of the wording,” which effectively utilize the convolutional layers to discern the finer details of the question.
  • In Case (b), the GRU model gives a completely wrong prediction of “picture” as it relies solely on simplistic word-level features. The ConvGRU(3) model improves by predicting “sitting”—a rationally deduced yet contextually incorrect response, revealing that 3-gram features effectively capture a person’s state or action in a specific location. The ConvGRU(2+3), by incorporating both 2-gram and 3-gram features, correctly comprehends the query’s intent and the subject’s context, leading to an accurate “reading” prediction. This case exemplifies how combining different n-grams can enrich linguistic feature understanding and enhance alignment with image features for more accurate predictions.
  • In both the counting task and the yes/no task, as shown in Case (c) and Case (d), ConvGRU(2+3) showcases best performance. The ability to understand relational context and specific details allows it to correctly count “3” boys on the field and accurately confirm the presence of avocado on the sandwich with a “yes”. These outcomes illustrate the effectiveness of integrating multi-gram features into GRUs.
@@ 5 Conclusion
This paper explored the effectiveness of complex versus simpler models for capturing textual features in Visual Question Answering (VQA) tasks, specifically addressing whether complex sequential models are optimal for this purpose. Our experiments demonstrated that the ConvGRU model, designed to capture local text features, marginally outperformed the baseline GRU model, while advanced models like Self-Attention and Transformer Encoders showed reduced effectiveness. The primary advantage of ConvGRU lies in its ability to balance model performance with efficiency.